# Analytical results for the Quantum Alternating Operator Ansatz with Grover Mixer

### v2

@@ III.1 Threshold curve problem
An open question in GM-Th-QAOA is about the behavior of the expectation value as a function of the threshold, fixing optimal $\bm{\beta}$ and $\bm{\gamma}$ for a given choice of $t$. In the present work, we call this function a threshold curve. It has been numerically observed that the threshold curve decreases monotonically up to a valley value and then increases monotonically [24]—recall that we are considering minimization problems. If that is always the case, a modified binary search is applicable to find the optimal threshold, representing an exponential improvement over the required linear search otherwise.
@@ III.2 Asymptotic tight bound on quantile
An alternative metric on the performance of GM-Th-QAOA over the expectation value and approximation ratio is the quantity $F_{X}(E_{r}(t))$, which corresponds to the quantile in which the expectation value of GM-Th-QAOA is associated. That metric has as a strong point the possibility of comparing the obtained result with the spectrum of distribution itself, which allows a fair comparison between different distributions and optimization problems. The immediate upper bound $F_{X}(E_{r}(t)_{opt})\leq F_{X}(\tau)$ can be obtained by applying the cdf to both sides of the inequality in Corollary 1. If we assume a continuous distribution, there is a $t$ in which $F_{X}(t)=\rho_{Th}(r)$ for all $r$ and then $F_{X}(\tau)=\rho_{Th}(r)$ for any $r$. That way,
@@ III.3 Upper bounds on the standard score
The corollary implies that $E_{r}(t)$ deviates from the mean proportionally to $\sigma$. Therefore, for a given $X$, it is straightforward to consider the negative of the standard score, i.e., a $C_{r}(t)$ such that $E_{r}(t)=\mu-C_{r}(t)\sigma$, as another performance metric. Note that the performance of GM-Th-QAOA, quantified by $C_{r}(t)$, is invariant over shifting the location of the distribution or changing its scale, being therefore dependent only on the statistical moments of higher order, such as skewness and kurtosis.
We can get the same conclusion about the expectation value of GM-QAOA with the analysis of Headley and Wilhelm [27]. From Equations (D15) and (D17) of Headley and Wilhelm [27] paper, denoting the characteristic function by $\varphi_{X}(\gamma)=\sum_{x\in R_{X}}f_{X}(x)e^{i\gamma x}$ and setting $B(\beta)=-1+e^{i\beta}$,
@@
Indeed, the above bound is not tight. The tight upper bound is established through the assistance of Lemma 1, which claims that $C^{Th}(r)$ is attained with a particular family of distribution: the two-point distributions. For such distributions, since the random variable $Z$ depends only on the ratio between the probability of the points, we can refer to our defined binary function without loss of generality.
@@
Proof:
$\begin{split}&\operatorname{E}[Y_{\leq T}]=\frac{G_{Y}(T)}{F_{Y}(T)},\ \operatorname{E}[Y_{\leq T}^{2}]=\frac{\sigma^{2}_{\leq T}}{F_{Y}(T)},\\ &\operatorname{E}[Y_{>T}]=-\frac{G_{Y}(T)}{1-F_{Y}(T)},\ \operatorname{E}[Y_{>T}^{2}]=\frac{\sigma^{2}_{>T}}{1-F_{Y}(T)},\end{split}$ (34)
Combining both bounds gives
$\frac{|G_{Y}(T)|}{\sigma}\leq\sqrt{F_{Y}(T)(1-F_{Y}(T))},$ (36)
and the equality is hit on the limit of Jensen’s inequality if and only if both $Y_{\leq T}$ and $Y_{>T}$ are degenerate distributions. $\square$
@@
For any number $r$ of layers on GM-Th-QAOA, provided that $R_{X}^{min}\neq 0$ and $|R_{X}^{min}|<\infty$,
@@ III.4 Combining the bounds on the standard score and quantile
In general, the optimal $C_{r}(t)$ asymptotically depends on $r$ as $C_{r}(t)=\Theta(|F^{-1}_{Z}(1/r^{2})|)$. If a distribution presents a cdf $F_{Z}(x)$ with exponential decay on $x\rightarrow-\infty$, the growth of $C_{r}(t)$ must be logarithm. It is the case of important distributions of literature, such as the normal, Laplace, gamma (reflected), and exponential (reflected) distributions. Therefore, for an optimization problem with a probability distribution that exhibits a tendency of exponential decay, the number of rounds to achieve a fixed approximation ratio must be exponentially larger than the tight bound of Corollary 3.
@@ IV Numerical experiments
We begin with the normal distribution. The study of that distribution is justified by the ubiquity generated by the central limit theorem. In combinatorial optimization context, empirically has been observed that the solution space of the Capacitated Vehicle Routing and Portfolio Optimization seems to be normally distributed [22, 33]. It would not be surprising if other optimization problems were normally distributed.
The choice of parameters $u$ and $s^{2}$ on normal distribution is irrelevant to the analysis since the random variable $Z$ is always the standard normal distribution. The results of QAOA are compared with a brute force algorithm known as classical random sampling (CRS) [22], which samples a given number of times of the distribution and takes the smallest one. As Bennett et al. [22], the equivalent computational effort of CRS to be used is $2r$. The minimum value of a set of samples is known as the first order statistic. The expected value of the first order statistic of the normal distribution is computed by using the Blom [42] asymptotic approximation,
@@ V General bounds on Grover-based QAOA
In Sec. III, we prove an asymptotic tight bound that implies GM-Th-QAOA has a quadratic speed up over the classical brute force approach. That raises the question of whether that bound is general for any Grover-based QAOA, whether GM-QAOA, or any potential new variation that can emerge. To answer that question, we develop a technique to get a general upper on Grover-based QAOA that consists of getting the maximum amplification of probability over any set of degenerate states. With that upper bound established, we explicitly construct the minimum expectation value within that constrained framework.
@@
To bound the maximum amplification of the probability over any set of degenerate states, we define $S_{T}$ as a set of elements on the spectrum of the Hamiltonian $H_{Q}$ with some fixed cost $x_{o}$. Suppose that $H_{Q}$ is built from an arbitrary problem Hamiltonian $H_{C}$. For a given $r$, our goal is to maximize the ratio between the probability of measuring a state on $S_{T}$ before and after the application of the QAOA operators optimizing the choices of the ratio $|S_{T}|/M$ and the probability distribution $f_{X_{q}}(x)$. The only restriction on the choice of the distribution $f_{X_{q}}(x)$ is sign the probability $|S_{T}|/M$ on value $x_{o}$. To get it, consider taking the expectation value on the final state of Grover-based QAOA of a third Hamiltonian $H_{max}$ that encodes $x_{o}$ to $|S_{T}|$ elements and 00 to the remainders, with ratio $\rho=|S_{T}|/M$. The probability of measuring an element of $S_{T}$ on the initial state is $\rho$, while after the application of QAOA operators is $E_{r}^{max}(S_{T},f_{X_{q}})/x_{o}$, where $E_{r}^{max}(S_{T},f_{X_{q}})$ denotes the expectation value of that configuration. We want to maximize the ratio between then, named $\eta_{r}(S_{T},f_{X_{q}})$, and use it bound to explicitly build the minimum expectation value on an arbitrary instance of some Grover-based QAOA by sequentially maximally amplifying the states in ascending order of cost until the sum of probabilities reaches $1$. As the amplitudes of degenerate states are equal, the amplification is in ascending order of the support of $X_{c}$.
@@
Note that the maximum amplification is $(2r+1)^{2}$, the exact amplification of Grover’s algorithm on the low-convergence regime. One can ask if the maximum amplification is on the low-convergence regime for any $r$. There is numerical evidence for that from Bennett and Wang [33] in the context of QWOA on the complete graph. Unfortunately, applying the individual bounds of $|B(\beta)|$ and $|\varphi_{X_{q}}(\gamma)|$ on general $r$ expression gives an exponential amplification of $9^{r}$. Moreover, due to the complexity of the expression, direct analytical treatment is unfeasible, necessitating indirect methods. Specifically, we demonstrate in Lemma 2, proved on Appendix B, that the maximum amplification is $(2r+1)^{2}$ by showing that the existence of a distribution that can achieve a larger amplification implies in an explicit algorithm for the unstructured search problem with a larger average probability than the bound of Hamann, Dunjko, and Wölk [36]. With Lemma 2, we can prove the lower bound on $E_{r}$ (i.e., a general upper bound on Grover-based QAOA performance), given by Theorem 5.
@@
For any number $r$ of layers on Grover-based QAOA with a set $S_{T}$ of ratio $\rho$, the amplification of the probability of measuring the elements of $S_{T}$ is bounded by
@@
where the tight bound is achieved with $\rho\rightarrow 0$ and $q(k)$ equal to the binary function up to a scale change.
@@
For any number $r$ of layers on Grover-based QAOA, the expectation value is bounded by
@@
where $\tau_{1}$ is the maximum element of the support of $X_{c}$ in which $F_{X_{c}}(t)\leq 1/(2r+1)^{2}$ and $\tau_{2}$ minimum element in which $F_{X_{c}}(t)>1/(2r+1)^{2}$. In particular, if $F_{X_{c}}(\tau_{1})=1/(2r+1)^{2}$, then $E_{r}\geq\operatorname{E}[{X_{c}}|{X_{c}}\leq\tau_{1}]$.
@@
and thus GM-Th-QAOA is, in the worst case, $\pi^{2}/4$ times worse than the maximum amplification bound in terms of the cdf. With the maximum amplification bound, we can also bound $C(r)$—and the number of rounds to achieve a fixed approximation ratio—obtaining the analogous of Theorem 4 and Corollary 3 for Grover-based QAOA, synthesized in Theorem 6.
@@
For any number $r$ of layers on Grover-based QAOA, $C(r)\leq 2\sqrt{r(r+1)}$ and, provided that $R_{X}^{min}\neq 0$ and $|R_{X}^{min}|<\infty$,
@@
Proof:
with $F_{X_{1}}(\tau_{2})=1/(2r+1)^{2}$. Now, we can apply the same proceeding of Lemma 1 on the random variables $X_{\leq\tau_{2}}$, the random variable $X_{1}$ given $X_{1}\leq\tau_{2}$; and $X_{\geq\tau_{2}}$, the random variable $X_{2}$ given $X_{2}\geq\tau_{2}$, to get the same conclusion. Then, take binary function on Eq. (53) gives $C(r)\leq 2\sqrt{r(r+1)}$. $\square$
@@
To finish, a direct comparison with Grover Adaptive Search follows directly from the bound on amplification of Lemma 2. Since the probability is bounded by $f_{X_{c}}(R_{X_{c}}^{min})(2r+1)^{2}$, finding an optimal solution for an optimization problem with Grover-based QAOA with probability of at least, for instance, $1/2$, needs $\Omega(1/\sqrt{f_{X_{c}}(R_{X_{c}}^{min})})$ rounds, an analogous result to the Theorem 1 of Durr and Hoyer [37].
@@
Figure 8: (a) Linear-log base $2$ plot of approximation ratio versus the number of layers considering the application of the maximum amplification bound for the graph $K_{50,50}$ on Max-Cut. As expected, the growth rate is logarithmic. The resolution used on the number of layers (due to the extremely high number of layers to get $\lambda=1$) is $\lceil 2^{x/100}\rceil$ for $x=0,1,\ldots,5000$. (b) The minimum number of layers required to maximum amplification bound achieves three different values of approximation ratio on the Max-Cut with the graph $K_{n,n}$ for $n=4,5,\ldots,100$. The considered values of $\lambda$ are $\lambda=1$, in which we calculate analytically that $r=\lceil 2^{0.5(|\mathcal{V}|-3)}\rceil$; $\lambda=16/17\approx 0.9411$, the $\lambda$ value in which Max-Cut becomes NP-hard [44, 45]; and the approximation ratio guaranteed by the classical Goemans-Williamson algorithm [46], given by $\lambda\approx 0.8786$. The scale is log-linear with base $2$. The value of $r$ in which the approximation ratio is achieved was efficiently found with a binary search. As predicted, the number of layers scales exponentially in all of them. (c) The same as (b), but with $n=4,5,\ldots,300$ and $\lambda=0.52$. As the expectation value of a uniform superposition gives $\lambda=0.5$, this approximation ratio is extremely low. However, even for such low performance, given a sufficient number of vertices, we observe the exponential dependence on the number of layers $r$.
V.1 Bounds on Max-Cut
One application of our bounds is the Max-Cut, a problem widely considered in QAOA literature [47]. The Max-Cut problem consists of finding a partition of the vertices of a given graph into two complementary subsets that maximize the number of edges between both subsets. Firstly, Theorem 6 improves by a constant factor the lower bound of order $\Omega(\sqrt{|\mathcal{E}|})$, where $|\mathcal{E}|$ is the number of edges of the graph, on bipartite graphs, given by Equation (36) of Benchasattabuse et al. [43] paper. In general, if the statistical quantity $(\mu-R_{X_{c}}^{min})/\sigma$ grows with the size of the instance to a given combinatorial optimization problem, we cannot achieve a fixed approximation ratio with a constant number of layers, a severe limitation for the NISQ context. However, at least in Max-Cut, the situation seems to be even worse, as we can see applying the bound of Eq. (54).
@@ V.1 Bounds on Max-Cut
Let us consider the complete bipartite graph $K_{n,n}$ with bipartition on the sets $V_{1}$ and $V_{2}$. Suppose a solution of Max-Cut as a partition on the sets $S_{1}$ and $S_{2}$ in which among the vertices of $S_{1}$, the number of vertices that belong to $V_{1}$ and $V_{2}$ are respectively $j$ and $k$. The size of the cut can be computed by discounting to the total number of edges $n^{2}$ of the whole graph, $jk+(n-j)(n-k)$, i.e., the number of edges induced by the union of both complete bipartite graphs $K_{j,k}$ and $K_{n-j,n-k}$ within the sets $S_{1}$ and $S_{2}$, respectively. Thus, with our definition of considering minimization problems and using the random variable $Y$ by subtraction of the mean $-n^{2}/2$, the solution space is composed by $\binom{n}{j}\binom{n}{k}$ solutions of cost $\frac{1}{2}(n-2j)(n-2k)$ for all $0\leq j,k\leq n$. Although we do not have the explicit distribution, the current characterization of the solution space is sufficient to see that it presents an exponential decay toward the optimal solution. As the distribution is symmetric, we consider, by simplicity, $j,k\leq\lfloor n/2\rfloor$. Fixing $j$ (or $k$), with a large enough $n$, the number of solutions grows exponentially with $k$ (or $j$) at the same time that the costs have a linear decrease. That combined behavior implies that a trend of a linear increase of the costs to lower values of $k$ and $j$ is accompanied by an exponential decay of their probabilities. Fig. 7 illustrate the decay of the distribution for the graph $K_{50,50}$ by showing the graphics of $f_{Y}(x)$ and $F_{Y}(x)$. The exponential decay, with the arguments of Sec. III.4 on the asymptotic limit, infers in a logarithmic increase of $C_{r}$ with the number of layers, and therefore, as $(\mu-R_{X_{c}}^{min})/\sigma$ grows with the square root on the number of edges for bipartite graphs, the number of layers to a achieves a fixed $\lambda$ must increase exponentially with the number of vertices/edges.

### v3

@@ III.1 Threshold curve problem
An open question in GM-Th-QAOA is about the behavior of the expectation value as a function of the threshold, fixing optimal $\bm{\beta}$ and $\bm{\gamma}$ for a given choice of $t$. In the present work, we call this function a threshold curve. It has been numerically observed that the threshold curve decreases monotonically up to a valley value and then increases monotonically [24]—recall that we are considering minimization problems. In the numerical experiments of Sec. IV, we show in Fig. 6 an example of that behavior using quantities for abscissa and ordinate in which the threshold curve is equivalent. If that property holds in general, a modified binary search is applicable to find the optimal threshold, representing an exponential improvement over the required linear search otherwise.
@@ III.2 Asymptotic tight bound on quantile
An alternative metric on the performance of GM-Th-QAOA over the expectation value and approximation ratio is the quantity $F_{X}(E_{r}(t))$, which corresponds to the quantile in which the expectation value of GM-Th-QAOA is associated. That metric has as a strong point the possibility of comparing the obtained result with the spectrum of the distribution itself, which allows a fair comparison between different distributions and optimization problems. The immediate upper bound $F_{X}(E_{r}(t)_{opt})\leq F_{X}(\tau)$ can be obtained by applying the cdf to both sides of the inequality in Corollary 1. If we assume a continuous distribution, there is a $t$ in which $F_{X}(t)=\rho_{Th}(r)$ for all $r$ and then $F_{X}(\tau)=\rho_{Th}(r)$ for any $r$. That way,
@@ III.3 Upper bounds on the standard score
The corollary implies that $E_{r}(t)$ deviates from the mean proportionally to $\sigma$. Therefore, for a given $X$, it is straightforward to consider the negative of the standard score, i.e., a $C_{r}(t)$ such that
$E_{r}(t)=\mu-C_{r}(t)\sigma,$ (32)
as another performance metric. Note that the performance of GM-Th-QAOA, quantified by $C_{r}(t)$, is invariant over shifting the location of the distribution or changing its scale, being therefore dependent only on the statistical moments of higher order, such as skewness and kurtosis.
We can get the same conclusion about the expectation value of GM-QAOA with the analysis of Headley and Wilhelm [28]. From Equations (D15) and (D17) of Headley and Wilhelm [28] paper, denoting the characteristic function by
$\varphi_{X}(\gamma)=\sum_{x\in R_{X}}f_{X}(x)e^{i\gamma x}$ (33)
and setting $B(\beta)=-1+e^{i\beta}$,
@@
Indeed, the above bound is not tight. The tight upper bound is established through the assistance of Lemma 1, proved in Appendix B, which claims that $C^{Th}(r)$ is attained with a particular family of distribution: the two-point distributions. For such distributions, since the random variable $Z$ depends only on the ratio between the probability of the points, we can refer to our defined binary function without loss of generality.
@@
For any number $r$ of layers in GM-Th-QAOA, provided that $R_{X}^{min}\neq 0$ and $|R_{X}^{min}|<\infty$,
@@ III.4 Combining the bounds on the standard score and quantile
In general, the optimal $C_{r}(t)$ asymptotically depends on $r$ as $C_{r}(t)=\Theta(|F^{-1}_{Z}(1/r^{2})|)$. If a distribution presents a cdf $F_{Z}(x)$ with exponential decay on $x\rightarrow-\infty$, the growth of $C_{r}(t)$ must be logarithmic. It is the case of important distributions of literature, such as the normal, Laplace, gamma (reflected), and exponential (reflected) distributions. Therefore, for an optimization problem with a probability distribution that exhibits a tendency of exponential decay, the number of rounds to achieve a fixed approximation ratio must be exponentially larger than the tight bound of Corollary 3.
@@ IV Numerical experiments
We begin with the normal distribution. The study of that distribution is justified by the ubiquity generated by the central limit theorem. Recall the normally distributed instances of the Capacitated Vehicle Routing and Portfolio Optimization problems [22, 34]. It would not be surprising if other optimization problems were normally distributed. The choice of parameters $u$ and $s^{2}$ on normal distribution is irrelevant to the analysis since the random variable $Z$ is always the standard normal distribution. The results of QAOA are compared with a brute force algorithm known as classical random sampling (CRS) [22], which samples a given number of times of the distribution and takes the smallest one. As Bennett et al. [22], the equivalent computational effort of CRS to be used is $2r$. The minimum value of a set of samples is known as the first order statistic. The expected value of the first order statistic of the normal distribution is computed by using the Blom [44] asymptotic approximation,
@@ V General bounds on Grover-based QAOA
In Sec. III, we prove an asymptotic tight bound that implies GM-Th-QAOA has a quadratic speed-up over the classical brute force approach. That raises the question of whether that bound is general for any Grover-based QAOA, whether GM-QAOA or any potential new variation that can emerge. To answer that question, we develop a technique to get a general upper bound on Grover-based QAOA that consists of getting the maximum amplification of probability over any set of degenerate states. With that upper bound established, we explicitly construct the minimum expectation value within that constrained framework.
@@
To bound the maximum amplification of the probability over any set of degenerate states, we define $S_{T}$ as a set of elements on the spectrum of the Hamiltonian $H_{Q}$ with some fixed cost $x_{o}$. Suppose that $H_{Q}$ is built from an arbitrary problem Hamiltonian $H_{C}$. For a given $r$, our goal is to maximize the ratio between the probability of measuring a state on $S_{T}$ before and after the application of the QAOA operators optimizing the choices of the ratio $|S_{T}|/M$ and the probability distribution $f_{X_{q}}(x)$. The only restriction on the choice of the distribution $f_{X_{q}}(x)$ is sign the probability $|S_{T}|/M$ on value $x_{o}$. To get it, consider taking the expectation value on the final state of Grover-based QAOA of a third Hamiltonian $H_{max}$ that encodes $x_{o}$ to $|S_{T}|$ elements and 00 to the remainders, with ratio $\rho=|S_{T}|/M$. The probability of measuring an element of $S_{T}$ on the initial state is $\rho$, while after the application of QAOA operators is $E_{r}^{max}(S_{T},f_{X_{q}})/x_{o}$, where $E_{r}^{max}(S_{T},f_{X_{q}})$ denotes the expectation value of that configuration. We want to maximize the ratio between then, named $\eta_{r}(S_{T},f_{X_{q}})$, and use it bound to explicitly build the minimum expectation value on an arbitrary instance of some Grover-based QAOA by sequentially maximally amplifying the states in ascending order of costs until the sum of probabilities reaches $1$. As the amplitudes of degenerate states are equal, the amplification is in ascending order of the support of $X_{c}$.
@@
Note that the maximum amplification is $(2r+1)^{2}$, the exact amplification of Grover’s algorithm on the low-convergence regime. One can ask if the maximum amplification is on the low-convergence regime for any $r$. There is numerical evidence for that from Bennett and Wang [34] in the context of QWOA on the complete graph. Unfortunately, applying the individual bounds of $|B(\beta)|$ and $|\varphi_{X_{q}}(\gamma)|$ on general $r$ expression gives an exponential amplification of $9^{r}$. Moreover, due to the complexity of the expression, direct analytical treatment is unfeasible, necessitating indirect methods. Specifically, we demonstrate in Lemma 2, proved on Appendix C, that the maximum amplification is $(2r+1)^{2}$ by showing that the existence of a distribution that can achieve a larger amplification implies an explicit algorithm for the unstructured search problem with a larger average probability than the bound of Hamann, Dunjko, and Wölk [37]. With Lemma 2, we can prove the lower bound on $E_{r}$ (i.e., a general upper bound on Grover-based QAOA performance), given by Theorem 5.
@@
For any number $r$ of layers in Grover-based QAOA with a set $S_{T}$ of ratio $\rho$, the amplification of the probability of measuring the elements of $S_{T}$ is bounded by
@@
where the tight bound is achieved considering the limit of $\rho\rightarrow 0$ and $q(k)$ equal to the binary function up to a scale change.
@@
For any number $r$ of layers in Grover-based QAOA, the expectation value is bounded by
@@
where $\tau_{1}$ is the maximum element of the support of $X_{c}$ in which $F_{X_{c}}(t)\leq 1/(2r+1)^{2}$ and $\tau_{2}$ is the minimum element in which $F_{X_{c}}(t)>1/(2r+1)^{2}$. In particular, if $F_{X_{c}}(\tau_{1})=1/(2r+1)^{2}$, then $E_{r}\geq\operatorname{E}[{X_{c}}|{X_{c}}\leq\tau_{1}]$.
@@
and thus GM-Th-QAOA is, in the worst case, $\pi^{2}/4$ times worse than the maximum amplification bound in terms of the cdf. With the maximum amplification bound, we can also bound $C(r)$—and the number of rounds to achieve a fixed approximation ratio—obtaining the analogous of Theorem 4 and Corollary 3 for Grover-based QAOA, synthesized in Theorem 6, that is proved in Appendix D.
@@
For any number $r$ of layers in Grover-based QAOA, $C(r)\leq 2\sqrt{r(r+1)}$ and, provided that $R_{X}^{min}\neq 0$ and $|R_{X}^{min}|<\infty$,
@@
Furthermore, a direct comparison with Grover Adaptive Search follows directly from the bound on amplification of Lemma 2. Since the probability is bounded by $f_{X_{c}}(R_{X_{c}}^{min})(2r+1)^{2}$, finding an optimal solution for an optimization problem with Grover-based QAOA with probability of at least, for instance, $1/2$, needs $\Omega(1/\sqrt{f_{X_{c}}(R_{X_{c}}^{min})})$ rounds, an analogous result to the Theorem 1 of Durr and Hoyer [38].
To conclude the analytical discussion, we emphasize that although the optimization metric considered for the observable throughout this paper is the expectation value, the result of Lemma 2, i.e., the bound on the amplification of the probability of sampled states, is an indication of quadratic gain that is independent of the particular metric of expectation value and may extend the key conclusion of our paper that the performance of Grover-based QAOA is limited to a quadratic speed-up over classical brute force for other metrics. In particular, an important weakness of expectation value as an optimization metric is that, except for the cases of instances that exhibit concentration of the sampled states to the average value on $M\rightarrow\infty$, such as Ref. [32], it is possible that we have a distribution of the sampled states in which the probability of measuring the best solutions is not adequately amplified, instead more amplifying the states with lower quality [46, 47, 48, 34]. Alternative optimization metrics were introduced in the literature to address this issue, including the Conditional Value-at-Risk (CVaR) [46], the Gibbs Objective Function [47], the probability of sampling states above certain quality [48], and even the probability of measuring optimal solutions [34]. However, the bound of Lemma 2 naturally remedies the discussed problem. In particular, as an example of what we did in the proof of Theorem 5 for the expectation value, we can assume the most favorable distribution of sampled states for combinatorial optimization, i.e., assuming the maximum quadratic amplifying the probability of the best states in ascending order of costs.
V.1 Bounds on Max-Cut problem
One application of our bounds is the Max-Cut, a problem widely considered in QAOA literature [49]. The Max-Cut problem consists of finding a partition of the vertices of a given graph into two complementary subsets that maximize the number of edges between both subsets. The classical Goemans-Williamson algorithm [50] provides the best-known approximation ratio for that problem with $\lambda\approx 0.8786$, which induces a point of comparison for the performance of quantum algorithms that aim to find approximate solutions, such as the QAOA. Under the assumption that the unique games conjecture holds, the approximation Goemans-Williamson gives the best approximation ratio guarantee for any polynomial-time algorithm on Max-Cut [51]. Without that unproven assumption, approximate Max-Cut with an approximation ratio $\lambda=16/17\approx 0.9411$ is proved to be NP-hard [52, 53]. In this sense, a widely known analytical result in QAOA literature is that for the particular class of $3$-regular graphs, QAOA with the original transverse field mixer and single layer guarantees an approximation ratio of $\lambda=0.6924$ [4]. That result was extended to $r=2$ and $r=3$, with approximation ratio guarantees of $\lambda=0.7559$ and $\lambda=0.7924$, respectively [30].
@@ V.1 Bounds on Max-Cut problem
Figure 8: (a) Linear-log base $2$ plot of approximation ratio versus the number of layers considering the application of the maximum amplification bound for the graph $K_{50,50}$ on Max-Cut. As expected, the growth rate is logarithmic. The resolution used on the number of layers (due to the extremely high number of layers to get $\lambda=1$) is $\lceil 2^{x/100}\rceil$ for $x=0,1,\ldots,5000$. (b) The minimum number of layers required to maximum amplification bound achieves three different values of approximation ratio on the Max-Cut with the graph $K_{n,n}$ for $n=4,5,\ldots,100$. The considered values of $\lambda$ are $\lambda=1$, in which we calculate analytically that $r=\lceil 2^{0.5(|\mathcal{V}|-3)}\rceil$; $\lambda=16/17$, the $\lambda$ value in which Max-Cut becomes NP-hard; and the approximation ratio guaranteed by the classical Goemans-Williamson algorithm, given by $\lambda\approx 0.8786$. The scale is log-linear with base $2$. The value of $r$ at which the approximation ratio is achieved was efficiently found with a binary search. As predicted, the number of layers scales exponentially in all of them. (c) The same as (b), but with $n=4,5,\ldots,300$ and $\lambda=0.52$. As the expectation value of a uniform superposition gives $\lambda=0.5$, this approximation ratio is extremely low. However, even for such low performance, given a sufficient number of vertices, we observe the exponential dependence on the number of layers $r$.
In the specific context of GM-QAOA, Benchasattabuse et al. [45], by extending a lower bound on the time of quantum annealing to the context of QAOA [54], prove that for the class of bipartite graphs, to achieve any constant approximation ratio guarantee, that variant requires a number of layers on the order of $\Omega(\sqrt{|\mathcal{E}|})$, where $|\mathcal{E}|$ is the number of edges of the graph. For the more general context of Grover-based QAOA, from the application of Theorem 6, we improve by a constant factor that lower bound—given by Equation (36) of Benchasattabuse et al. [45] paper—with
$r\geq\frac{2\lambda-1}{2\sqrt{1+1/r}}\sqrt{|\mathcal{E}|}.$ (54)
In other words, for any choice of approximation ratio $\lambda$, the number of layers for the result of the algorithm reaches $\lambda$ must grow at least in order of $\sqrt{\mathcal{E}}$. Therefore, we cannot guarantee any fixed approximation ratio with a constant number of layers, a severe limitation for the NISQ devices, which require shallow depth. In general, from Theorem 6, this problem will appear if the statistical quantity $(\mu-R_{X_{c}}^{min})/\sigma$ grows with the size of the instance to a given combinatorial optimization problem. However, at least in Max-Cut, the situation seems to be even worse, as we can see applying the bound of Eq. (53).
@@
Let us consider the complete bipartite graph $K_{n,n}$ with bipartition on the sets $V_{1}$ and $V_{2}$. Suppose a solution of Max-Cut as a partition on the sets $S_{1}$ and $S_{2}$ in which among the vertices of $S_{1}$, the number of vertices that belong to $V_{1}$ and $V_{2}$ are respectively $j$ and $k$. The size of the cut can be computed by discounting to the total number of edges $n^{2}$ of the whole graph, $jk+(n-j)(n-k)$, i.e., the number of edges induced by the union of both complete bipartite graphs $K_{j,k}$ and $K_{n-j,n-k}$ within the sets $S_{1}$ and $S_{2}$, respectively. Thus, with our definition of considering minimization problems and using the random variable $Y$ by subtraction of the mean $-n^{2}/2$, the solution space is composed by $\binom{n}{j}\binom{n}{k}$ solutions of cost $\frac{1}{2}(n-2j)(n-2k)$ for all $0\leq j,k\leq n$. Although we do not have the explicit distribution, the current characterization of the solution space is sufficient to see that it presents an exponential decay toward the optimal solution. As the distribution is symmetric, we consider, by simplicity, $j,k\leq\lfloor n/2\rfloor$. Fixing $j$ (or $k$), with a large enough $n$, the number of solutions grows exponentially with $k$ (or $j$) at the same time that the costs have a linear decrease. That combined behavior implies that a trend of a linear increase of the costs to lower values of $k$ and $j$ is accompanied by an exponential decay of their probabilities. Fig. 7 illustrate the decay of the distribution for the graph $K_{50,50}$ by showing the graphics of $f_{Y}(x)$ and $F_{Y}(x)$. The exponential decay, with the arguments of Sec. III.4 on the asymptotic limit, infers in a logarithmic increase of $C_{r}$ with the number of layers, and therefore, as $(\mu-R_{X_{c}}^{min})/\sigma$ grows with the square root on the number of edges for bipartite graphs, the number of layers to achieve guaranteed in terms of $\lambda$ must increase exponentially with the number of vertices/edges.
That result points out a limitation of Grover-based QAOA that is serious beyond the NISQ context since the exponential increase in the number of rounds implies that we cannot obtain a polynomial-time algorithm that guarantees any constant approximation ratio. In the context of comparative literature between QAOA mixers, our findings for QAOA with Grover mixer contrast with the aforementioned result that the QAOA with transverse field mixer guarantees a constant approximation ratio for $3$-regular graphs even with a single round.